# Caffarelli’s work on elliptic free boundary problems

### v1

@@ 1. Introduction
The claim that “elliptic free boundary problems form a field” is in great part due to the revolutionary ideas injected by Luis since his very first papers on the topic, cf. [7], and even giving an account of all the research produced by him on the subject is a demanding task. Making a somewhat limited choice of what to explain to the reader seemed unavoidable to me. On the other hand I will try to compensate such arbitrary and limited choice by keeping this article as nontechnical as possible, in the hope that it can be understood by a general audience of mathematicians. In order to follow of the discussion in this note a basic knowledge of measure theory, of functional analysis, and of very elementary facts about harmonic functions will suffice. In fact I will list immediately all the facts about harmonic functions which I will take for granted:
@@ 1.1. First roadmap
  • In Section 3 I will explain why the work of Luis is so groundbreaking: let me anticipate that prior to his papers almost nothing was known on the “regularity of the free boundary” of the solutions. With his works not only Luis produced theorems which lifted our almost empty knowledge to a very satisfactory level, but he also paved the way to a variety of later developments in the field. I will argue that there is an underlying philosophy to his approach to regularity which is quite general and powerful, and which it is in fact quite neatly explained in other references, however all the ones I know are written for researchers in PDE.
@@
  • Section 9 will serve as a teaser for further readings. I will mention how similar regularity theorems were proved by Luis and some of his collaborators in the context of the other three problems and I will point out to further readings in the area.
@@ 2.2. The thin obstacle
In the thin obstacle we again wish to minimize an elliptic energy (which for simplicity we assume again to be the Dirichlet energy) over a class of functions which this time are assumed to satisfy a boundary condition of Dirichlet type $u=g$ on some portion of the boundary and an inequality of the form $u\geq\varphi$ on the remaining portion. A model situation is the following:
@@ 2.3. The Bernoulli one-phase problem
is lower semicontinuous. The energy $E(u)$ is thus lower semicontinuous as well and the existence of a minimizer can be concluded from the direct methods.
@@ 3. Regularity of the free boundary: general themes
In all the four examples of the previous section deriving the existence of a minimizer of the corresponding variational problem is pretty straightforward (while, of course, we are taking advantage of a long history of functional analytic tools developed in the past decades). However the resulting minimizer may be far from the description given at the beginning of the section of a solution of a “free boundary problem”.
@@
Some results showing the expected regularity for the function $u$ when regarded as a function on the whole domain $D$ existed before Luis’ work, for instance the $C^{1,1}$ regularity of solutions to the obstacle problem was indeed proved by Frehse, cf. [10]. For the other three problems mentioned in the previous sections later works of Luis (some with coauthors) got the expected regularity, cf. [4, 1, 2].
@@
Prior to Luis’ first works in the area (i.e. essentially prior to [7]), the overall results in gaining some form of regularity for $\partial\Omega$ were almost inexistent: the ones I am ware of concern the first two of the four examples in Section 2: (notable) works of Hans Lewy, cf. [13, 12], in the cases of the obstacle and thin obstacle problems proved regularity for the free boundary in the $2$-dimensional setting. There were instead quite general works showing that some initial regularity of $\partial\Omega$ could be bootstrapped to higher one (cf. the work [11]).
The groundbreaking paper [7] not only gave a satisfactory answer to the boundary regularity of $\partial\Omega$ for the obstacle problem, but it also outlined a general strategy to attack the question which was later implemented by Luis himself and coauthors (e.g. in the other three examples of the previous section) and by other mathematicians in several other settings. The work outlined a general philosophy, which I will try to explain next.
@@
At this point a pivotal and very much desired property is to infer some restriction about the type of blow-ups which we might encounter. The reader will notice that, by the very way we got to these blow-ups, they are presumably “global” solutions of the corresponding variational/PDE problem, i.e. they are defined on the whole space and their restrictions to compact sets are minimizers of the same variational problem. We might then expect some “Liouville-type” theorem which classifies these objects. However, Liouville type theorems are pretty hard to get. In the various examples given in this note, some extra help come from additional information that can be inferred from the blow-up procedure. In the case of the obstacle problem this is for instance a powerful convexity estimate. In other cases an underlying monotonicity formula gives, for the blow-up, the constancy of an additional quantity [^3] Often these tools can indeed be used to infer a Liouville type theorem without knowing that the objects classified came as a “blow-up” limit, but we will disregard this aspect.. Notably, an interesting monotonicity formula was discovered later for the obstacle problem as well by Weiss (see [9] and the references therein), though it is not powerful enough to give, alone, a complete classification of blow-ups.
@@
Before getting further, a remark is in order. The pattern discovered by Luis in free boundary problems share many similarities with several other fundamental discoveries of similar nature in the calculus of variations, analysis of PDEs, and geometric analysis. A pioneering example, in fact probably the very first example in the literature, is given by the works of De Giorgi, Reifenberg, Almgren, and Allard in the context of minimal surfaces. In various articles and surveys Luis himself acknowledges the profound impact that the latter literature, and especially De Giorgi’s works in elliptic regularity theory, has had on him. However, reading the original paper [7] I came to the conclusion that the first work of Luis on the obstacle problem has a number of fundamental differences with the approach of the latter authors to regularity, and that the (fruitful!) connection with the minimal surface theory came rather in the later works.
@@ 3.1. Second roadmap
As claimed in the introduction, in the next five sections I will try to give, in as simple terms as possible, some details on the implementation of the above plan in the model case of the obstacle problem, following essentially [7], but taking very much advantage of the beautiful survey [9]. In Section 4 I will give the very first elementary remark, which will serve as starting point of our discussion, while in Section 5 I will discuss Frehse’s $C^{1,1}$ regularity theorem, which we will interpret as suitable bound from above on the growth of $v=u-\varphi$ as we get away from the free boundary. These sections describe essentially the state of the art of the regularity theory for the obstacle problem when [7] appeared, with the notable exception of the two-dimensional case. With the remaining three sections we enter into the fundamental contributions of Luis’ paper. Section 6 will discuss the matching lower bound, introduce the rescalings and their limits (the blow-ups). Section 7 will illustrate the fundamental convexity estimate which triggers the “classification” of the blow-ups and underlies the partitioning of the free boundary in a regular and singular part. Section 8 will finally outline the ideas behind the $\varepsilon$-regularity statement. In both these sections we will see the masterful use of the Harnack estimate for harmonic functions (and more in general for solution of elliptic PDEs) which underlie Luis’ regularity theory in [7].
@@ 5.1. Growth bound
Summig the last two inequalities we reach
@@ 5.2. Interior $C^{1,1}$ regularity
We recall here a simple effect of the mean-value property for harmonic functions, namely the estimate
@@
Cosider thus a point $x\in\Omega$ closer to $D\cap\partial\Omega$ then to $\partial D$, and set $r:=\operatorname{dist}(x,\partial\Omega)$. The function $y\mapsto w(y)=u(y)-\varphi(x)-\nabla\varphi(x)\cdot(y-x)$ is harmonic in $B_{r}(x)$ and we can apply the above estimate with $k=2$ to infer
$|D^{2}u(x)|=|D^{2}w(x)|\leq Cr^{-2}\sup_{B_{r}(x)}w\,.$
@@
Therefore both $v$ and $\nabla v$ extend continuously to zero from $\Omega$ to $D$, thus proving the continuous differentiability of $v$. For the Lipschitz regularity of $\nabla v$ one has to work slightly more, but essentially it is a matter of integrating the bound on the second derivative on the right paths.
@@ 6. Lower bound on the growth and blow-ups
What explained so far was known at the time the paper [7] appeared: this and the next two sections really enter in the main new ideas introduced by Luis. From now on the goal is really to study the regularity of the “free boundary” $\partial\{u>\varphi\}=\partial\{v>0\}$, but before starting let us observe that, without introducing some more assumptions, we cannot expect any regularity for it.
@@ 6.2. Blow-ups
Recall next that we have Frehse’s $C^{1,1}$ estimate. The latter is essentially an estimate on the second derivative: it says that, even though not continuous, it is bounded. The rescaling above preserves however the size of the second derivative, i.e. $D^{2}v_{r}(x)=D^{2}v(rx)$. Thus the Liphscitz constant of $\nabla v_{r}$ is uniformly bounded. Moreover, given that $\nabla v_{r}(0)=0$ and $v_{r}(0)=0$ we immediately infer that, on every fixed ball $B_{R}(0)\subset\mathbb{R}^{n}$, both $\|v_{r}\|_{C^{0}}$ and $\|\nabla v_{r}\|_{C^{0}}$ enjoy uniform bounds (the first with a quadratic dependence on $R$ and the second with a linear dependence). But then by the Ascoli-Arzelá theorem, for every sequence $r_{k}\downarrow 0$ there is a subsequence, not relabeled, such that $v_{r_{k}}$ converges uniformly on compact subsets of $\mathbb{R}^{2}$ to some function. It is also not difficult to see that the limit $v_{0}$ is as well $C^{1,1}$ and that $\nabla v_{0}(0)=0$ and $v_{0}(0)=0$. $v_{0}$ is what we will call “a blow-up” at the point $x_{0}\in\partial\Omega$ [^5] For general $x_{0}\in\partial\Omega$ the rescaled functions will take the form $v_{r}(x)=r^{-2}v(x_{0}+rx)$. This essentially amounts to first translate the solution so that $x_{0}$ is the origin, and then use the formula (6.3)..
@@ 7.3. Proof of the convexity estimate
Assume that $v(x_{0})>0$ and that $\partial B_{r}(x_{0})$ touches $\partial\Omega$ at a point $y\in D$, while $B_{r}(x_{0})\subset\Omega$. Then for every unit vector $e$ and for any $\delta<\frac{1}{2}$ there is a point $x_{1}$ in $B_{(1-\delta/2)r}(x_{0})$ at which we have the estimate
$D_{ee}v(x_{1})\geq-C\sqrt{\delta}$
@@
$0\leq v(x_{1})\leq v(\delta x_{0})+c_{0}\sqrt{\delta}r|D_{e}v(\delta(x_{0}))|+\tfrac{1}{2}c_{0}^{2}\delta r^{2}\sup_{\sigma_{1}}D_{ee}v\,.$
@@
$\tfrac{1}{2}c_{0}^{2}\delta r^{2}\sup_{\sigma}D_{ee}v\geq-\tfrac{1}{2}C_{0}(\delta r)^{2}-C_{0}\delta^{\frac{3}{2}}r^{2}\,.$
Since $\sigma\subset B_{(1-\delta/2)r}(x_{0})$ the claim of the lemma readily follows. ∎
@@
when $x$ is at distance at most $\frac{r}{2}$ from $D\cap\partial\Omega$. If $f(M_{0})$ were negative or 00 we would be rather happy (because we would immediately conquer $D_{ee}v\geq 0$ in a neighborhood of $D\cap\partial\Omega$ and away from $\partial D$). Otherwise we can keep iterating, i.e. at distance $2^{-k}r_{0}$ we will $-f^{k}(M_{0})$ as lower bound, where $f^{k}=f\circ\ldots\circ f$ is the $k$-th iterate of the map $f$. Observe that, because we are assuming $f(M_{0})>0$, we have immediately $0<f(M_{0})<M_{0}$, hence also $0<f^{2}(M_{0})<f(M_{0})$, and so on: $f^{k}(M_{0})$ stays always positive. It is now pretty simple to argue that the monotone sequence $M_{k}=f^{k}(M_{0})$ must in fact converge to 00, which means that our lower bound on $D_{ee}v$ will approach zero as we get closer and closer to $D\cap\partial\Omega$ (but staying away from $\partial D$!). Indeed if it were $M_{\infty}=\inf_{k}M_{k}=\lim_{k}M_{k}>0$, then setting
@@
Assume $f$ is a nonnegative harmonic function on $B_{R}(x_{0})\subset\mathbb{R}^{n}$ and let $x_{1}$ be a point in $B_{r}(x_{0})$ with $r<R$. Then
$\frac{1+(\tfrac{r}{R})^{n-1}}{1-\tfrac{r}{R}}\geq f(x_{0})\geq\frac{1-(\tfrac{r}{R})^{n-1}}{1+\tfrac{r}{R}}f(x_{1})\,.$
Now, to the proof of (I). Fix $r$ and $M$ as in the statement and let $x_{0}$ be a point at distance at most $\frac{r}{2}$ from $\partial\Omega$. In fact let $r^{\prime}$ be the largest radius centered at $x_{0}$ over which $v$ is positive, so that $\partial B_{r^{\prime}}(x_{0})$ touches $D\cap\partial\Omega$. Observe that $D_{ee}v$ is an harmonic function in $B_{r^{\prime}}(x_{0})$ (since $\Delta v\equiv 1$ in it, any partial derivative of $v$ is harmonic!). Moreover $D_{ee}v+M$ is positive by assumption (because $B_{r^{\prime}}(x_{0})$ is contained in the $r$-neighborhood of $\partial\Omega$). We apply now the Harnack inequality between $x_{0}$ and the point $x_{1}$ found in Lemma 7.4, where for the moment $\delta$ is not yet chosen. Then we find

### v2

@@ 1. Introduction
The claim that “elliptic free boundary problems form a field” is in great part due to the revolutionary ideas injected by Luis since his very first papers on the topic, cf. [7], and even giving an account of all the research produced by him on the subject is a demanding task. Making a somewhat limited choice of what to explain to the reader seemed unavoidable to me. On the other hand I will try to compensate such arbitrary and limited choice by keeping this article as nontechnical as possible, in the hope that it can be understood by a general audience of mathematicians. In order to follow the discussions in this note a basic knowledge of measure theory, of functional analysis, and of very elementary facts about harmonic functions will suffice. In fact I will list immediately all the facts about harmonic functions which I will take for granted:
@@ 1.1. First roadmap
  • In Section 3 I will explain why the work of Luis is so groundbreaking: let me anticipate that prior to his papers almost nothing was known on the “regularity of the free boundary” of the solutions. With his works not only Luis produced theorems which lifted our almost empty knowledge to a very satisfactory level, but he also paved the way to a variety of later developments in the field. I will argue that there is an underlying philosophy to his approach to regularity which is quite general and powerful, and which is in fact quite neatly explained in other references, however all the ones I know are written for researchers in PDE.
@@
  • Section 9 will serve as a teaser for further readings. I will mention how similar regularity theorems were proved by Luis and some of his collaborators in the context of the other three problems and I will point out further readings in the area.
@@ 2.2. The thin obstacle
In the thin obstacle problem we again wish to minimize an elliptic energy (which for simplicity we assume again to be the Dirichlet energy) over a class of functions which this time are assumed to satisfy a boundary condition of Dirichlet type $u=g$ on some portion of the boundary and an inequality of the form $u\geq\varphi$ on the remaining portion. A model situation is the following:
@@ 2.3. The Bernoulli one-phase problem
is lower semicontinuous. The energy $E(u)$ is thus lower semicontinuous as well and the existence of a minimizer can be concluded from the direct methods in the calculus of variations.
@@ 3. Regularity of the free boundary: general themes
In all four examples of the previous section, deriving the existence of a minimizer of the corresponding variational problem is pretty straightforward (while, of course, we are taking advantage of a long history of functional analytic tools developed in the past decades). However the resulting minimizer may be far from the description given at the beginning of the section of a solution of a “free boundary problem”.
@@
Some results showing the expected regularity for the function $u$ when regarded as a function on the whole domain $D$ existed before Luis’ work, for instance the $C^{1,1}$ regularity of solutions to the obstacle problem was indeed proved by Frehse, cf. [10]. For the other three problems mentioned in the previous sections later works of Luis (some with co-authors) got the expected regularity, cf. [4, 1, 2].
@@
Prior to Luis’ first works in the area (i.e. essentially prior to [7]), the overall results in gaining some form of regularity for $\partial\Omega$ were almost inexistent: the ones I am aware of concern the first two of the four examples in Section 2: (notable) works of Hans Lewy, cf. [13, 12], in the cases of the obstacle and thin obstacle problems proved regularity for the free boundary in the $2$-dimensional setting. There were instead quite general works showing that some initial regularity of $\partial\Omega$ could be bootstrapped to higher one (cf. the work [11]).
The groundbreaking paper [7] not only gave a satisfactory answer to the boundary regularity of $\partial\Omega$ for the obstacle problem, but it also outlined a general strategy to attack the question which was later implemented by Luis himself and co-authors (e.g. in the other three examples of the previous section) and by other mathematicians in several other settings. The work outlined a general philosophy, which I will try to explain next.
@@
At this point a pivotal and very much desired property is to infer some restriction about the type of blow-ups which we might encounter. The reader will notice that, by the very way we got to these blow-ups, they are presumably “global” solutions of the corresponding variational/PDE problem, i.e. they are defined on the whole space and their restrictions to compact sets are minimizers of the same variational problem. We might then expect some “Liouville-type” theorem which classifies these objects. However, Liouville type theorems are pretty hard to get. In the various examples given in this note, some extra help comes from additional information that can be inferred from the blow-up procedure. In the case of the obstacle problem this is for instance a powerful convexity estimate. In other cases an underlying monotonicity formula gives, for the blow-up, the constancy of an additional quantity [^3] Often these tools can indeed be used to infer a Liouville type theorem without knowing that the objects classified came as a “blow-up” limit, but we will disregard this aspect.. Notably, an interesting monotonicity formula was discovered later for the obstacle problem as well by Weiss (see [9] and the references therein), though it is not powerful enough to give, alone, a complete classification of blow-ups.
@@
Before getting further, a remark is in order. The pattern discovered by Luis in free boundary problems shares many similarities with several other fundamental discoveries of similar nature in the calculus of variations, analysis of PDEs, and geometric analysis. A pioneering example, in fact probably the very first example in the literature, is given by the works of De Giorgi, Reifenberg, Almgren, and Allard in the context of minimal surfaces. In various articles and surveys Luis himself acknowledges the profound impact that the latter literature, and especially De Giorgi’s works in elliptic regularity theory, has had on him. However, reading the original paper [7] I came to the conclusion that the first work of Luis on the obstacle problem has a number of fundamental differences with the approach of the latter authors to regularity, and that the (fruitful!) connection with the minimal surface theory came rather in the later works.
@@ 3.1. Second roadmap
As claimed in the introduction, in the next five sections I will try to give, in as simple terms as possible, some details on the implementation of the above plan in the model case of the obstacle problem, following essentially [7], but taking very much advantage of the beautiful survey [9]. In Section 4 I will give the very first elementary remark, which will serve as starting point of our discussion, while in Section 5 I will discuss Frehse’s $C^{1,1}$ regularity theorem, which we will interpret as a suitable bound from above on the growth of $v=u-\varphi$ as we get away from the free boundary. These sections describe essentially the state of the art of the regularity theory for the obstacle problem when [7] appeared, with the notable exception of the two-dimensional case. With the remaining three sections we enter into the fundamental contributions of Luis’ paper. Section 6 will discuss the matching lower bound, introduce the rescalings and their limits (the blow-ups). Section 7 will illustrate the fundamental convexity estimate which triggers the “classification” of the blow-ups and underlies the partitioning of the free boundary in a regular and singular part. Section 8 will finally outline the ideas behind the $\varepsilon$-regularity statement. In both these sections we will see the masterful use of the Harnack estimate for harmonic functions (and more in general for solution of elliptic PDEs) which underlie Luis’ regularity theory in [7].
@@ 5.1. Growth bound
Summing the last two inequalities we reach
@@ 5.2. Interior $C^{1,1}$ regularity
We recall here a simple consequence of the mean-value property for harmonic functions, namely the estimate
@@
Cosider thus a point $x\in\Omega$ closer to $D\cap\partial\Omega$ than to $\partial D$, and set $r:=\operatorname{dist}(x,\partial\Omega)$. The function $y\mapsto w(y)=u(y)-\varphi(x)-\nabla\varphi(x)\cdot(y-x)$ is harmonic in $B_{r}(x)$ and we can apply the above estimate with $k=2$ to infer
$|D^{2}u(x)|=|D^{2}w(x)|\leq Cr^{-2}\sup_{B_{r}(x)}|w|\,.$
@@
Therefore both $v$ and $\nabla v$ extend continuously to zero from $\Omega$ to $D$, thus proving the continuous differentiability of $v$. For the Lipschitz regularity of $\nabla v$ one has to work slightly more, but essentially it is a matter of integrating the bound for the second derivatives on the right paths.
@@ 6. Lower bound on the growth and blow-ups
What was explained so far was known at the time the paper [7] appeared: this and the next two sections really enter in the main new ideas introduced by Luis. From now on the goal is to study the regularity of the “free boundary” $\partial\{u>\varphi\}=\partial\{v>0\}$, but before starting let us observe that, without introducing some more assumptions, we cannot expect any regularity for it.
@@ 6.2. Blow-ups
Recall next that we have Frehse’s $C^{1,1}$ estimate. The latter is essentially an estimate on the second derivative: it says that, even though not continuous, it is bounded. The rescaling above preserves however the size of the second derivative, i.e. $D^{2}v_{r}(x)=D^{2}v(rx)$. Thus the Liphscitz constant of $\nabla v_{r}$ is uniformly bounded. Moreover, given that $\nabla v_{r}(0)=0$ and $v_{r}(0)=0$ we immediately infer that, on every fixed ball $B_{R}(0)\subset\mathbb{R}^{n}$, both $\|v_{r}\|_{C^{0}(B_{R}(0))}$ and $\|\nabla v_{r}\|_{C^{0}(B_{R}(0))}$ enjoy uniform bounds (the first with a quadratic dependence on $R$ and the second with a linear dependence). But then by the Ascoli-Arzelá theorem, for every sequence $r_{k}\downarrow 0$ there is a subsequence, not relabeled, such that $v_{r_{k}}$ converges uniformly on compact subsets of $\mathbb{R}^{2}$ to some function. It is also not difficult to see that the limit $v_{0}$ is as well $C^{1,1}$ and that $\nabla v_{0}(0)=0$ and $v_{0}(0)=0$. $v_{0}$ is what we will call “a blow-up” at the point $x_{0}\in\partial\Omega$ [^5] For general $x_{0}\in\partial\Omega$ the rescaled functions will take the form $v_{r}(x)=r^{-2}v(x_{0}+rx)$. This essentially amounts to first translate the solution so that $x_{0}$ is the origin, and then use the formula (6.3)..
@@ 7.3. Proof of the convexity estimate
Assume that $v(x_{0})>0$ and that $\partial B_{r}(x_{0})$ touches $\partial\Omega$ at a point $y\in D$, while $B_{r}(x_{0})\subset\Omega$. Then for every unit vector $e$ and for any $\delta<\frac{1}{2}$ there is a point $\bar{x}_{1}$ in $B_{(1-\delta/2)r}(x_{0})$ at which we have the estimate
$D_{ee}v(\bar{x}_{1})\geq-C\sqrt{\delta}$
@@
$0\leq v(x_{1})\leq v(\delta x_{0})+c_{0}\sqrt{\delta}r|D_{e}v(\delta x_{0})|+\tfrac{1}{2}c_{0}^{2}\delta r^{2}\sup_{\sigma_{1}}D_{ee}v\,.$
@@
$\tfrac{1}{2}c_{0}^{2}\delta r^{2}\sup_{\sigma_{1}}D_{ee}v\geq-\tfrac{1}{2}C_{0}(\delta r)^{2}-C_{0}\delta^{\frac{3}{2}}r^{2}\,.$
Since $\sigma_{1}\subset B_{(1-\delta/2)r}(x_{0})$ the claim of the lemma readily follows. ∎
@@
when $x$ is at distance at most $\frac{r}{2}$ from $D\cap\partial\Omega$. If $f(M_{0})$ were negative or 00 we would be rather happy (because we would immediately conquer $D_{ee}v\geq 0$ in a neighborhood of $D\cap\partial\Omega$ and away from $\partial D$). Otherwise we can keep iterating, i.e. at distance $2^{-k}r_{0}$ we will have $-f^{k}(M_{0})$ as lower bound, where $f^{k}=f\circ\ldots\circ f$ is the $k$-th iterate of the map $f$. Observe that, because we are assuming $f(M_{0})>0$, we have immediately $0<f(M_{0})<M_{0}$, hence also $0<f^{2}(M_{0})<f(M_{0})$, and so on: $f^{k}(M_{0})$ stays always positive. It is now pretty simple to argue that the monotone sequence $M_{k}=f^{k}(M_{0})$ must in fact converge to 00, which means that our lower bound on $D_{ee}v$ will approach zero as we get closer and closer to $D\cap\partial\Omega$ (but staying away from $\partial D$!). Indeed if it were $M_{\infty}=\inf_{k}M_{k}=\lim_{k}M_{k}>0$, then setting
@@
Assume $f$ is a nonnegative harmonic function on $B_{R}(x_{0})\subset\mathbb{R}^{n}$ and let $\bar{x}_{1}$ be a point in $B_{r}(x_{0})$ with $r<R$. Then
$f(\bar{x}_{1})\frac{1+(\tfrac{r}{R})^{n-1}}{1-\tfrac{r}{R}}\geq f(x_{0})\geq\frac{1-(\tfrac{r}{R})^{n-1}}{1+\tfrac{r}{R}}f(\bar{x}_{1})\,.$
Now, to the proof of (I). Fix $r$ and $M$ as in the statement and let $x_{0}$ be a point at distance at most $\frac{r}{2}$ from $\partial\Omega$. In fact let $r^{\prime}$ be the largest radius centered at $x_{0}$ over which $v$ is positive, so that $\partial B_{r^{\prime}}(x_{0})$ touches $D\cap\partial\Omega$. Observe that $D_{ee}v$ is an harmonic function in $B_{r^{\prime}}(x_{0})$ (since $\Delta v\equiv 1$ in it, any partial derivative of $v$ is harmonic!). Moreover $D_{ee}v+M$ is positive by assumption (because $B_{r^{\prime}}(x_{0})$ is contained in the $r$-neighborhood of $\partial\Omega$). We apply now the Harnack inequality between $x_{0}$ and the point $\bar{x}_{1}$ found in Lemma 7.4, where for the moment $\delta$ is not yet chosen. Then we find